When solid CaCO3 is heated, it decomposes to give solid CaO and CO2 gas. A volume of 360 mL of gas is collected over water at a total pressure of 730 mmHg and 16 ∘C. The vapor pressure of water at 16 ∘C is 14 mmHg. CaCO3(s)→CaO(s)+CO2(g)

a) What was the partial pressure of the CO2 gas? Express your answer with the appropriate units.

b) How many moles of CO2 gas were in the CO2 gas sample? Express your answer with the appropriate units.
Answer #1

The total pressure of 730 mm of hg is the partial pressure of carbondioxide gas because on decomposition of calcium carbonate the only gas produced is carbon dioxide.

a) partial pressure of carbon dioxide gas is 730mm of hg.

b) 760 mm of hg contains = 22400 ml of carbondioxide

730 mm of hg contains = 22400/760 multiply by 730= 21515.8 ml

So 21515.8 ml contains 1 mole at 730 mm of hg

360 ml of gas contains = 360/21515 = 0.016 mole of carbon dioxide.